# Unleashing Reasoning Capability of LLMs via Scalable Question Synthesis from Scratch

### v1

@@ 3.2 Main Results
ScaleQuest significantly outperforms others
Table 1 presents the results. ScaleQuest significantly outperforms previous synthetic methods, with average performance improvements ranging from 5.6% to 11.5% over the prior state-of-the-art (SoTA) on both general base models and math-specialized foundation models. Qwen2-Math-7B-ScaleQuest achieved a zero-shot pass@1 accuracy of 73.4 on the MATH benchmark, matching the performance of GPT-4-Turbo. For out-of-domain tasks, Qwen2-Math-7B-ScaleQuest outperformed its teacher model, Qwen2-Math-7B-Instruct, with scores of 89.7 on the GSM8K benchmark, 73.4 on the MATH benchmark, and 38.5 on the Olympiad benchmark. It’s important to highlight that Qwen2-Math-7B-Instruct has undergone Group Relative Policy Optimization (GRPO) (Shao et al., 2024), utilizing the powerful reward model Qwen2-Math-RM-72B (Yang et al., 2024a), while our model is only an instruction tuning version. To ensure a fair comparison with other baselines, we have only applied supervised fine-tuning (SFT) in this work, leaving the preference tuning process for future work.
@@
We also explored the scalability of our dataset. We used our constructed dataset along with publicly available datasets, including MetaMath (Yu et al., 2023a), DART-Math (Tong et al., 2024), and Numina-Math (Li et al., 2024c). We trained the model using Llama3-8B and observed how its performance scaled with increasing data size. The results are presented in Figure 1. For the in-domain evaluation (MATH), our method demonstrates high data efficiency, achieving superior results with the same amount of data. In out-of-domain evaluations (Olympiad Bench), it also shows strong scalability, continuing to improve even as other datasets reach their limits. A limited question set leads to constrained improvements in model performance, as demonstrated by the results of DART-Math, which relies on a small number of questions and generates numerous correct answers through rejection sampling. Limited questions face a scalability ceiling, as the lack of diversity in the question set restricts further performance growth. Our results further demonstrate that diverse questions support sustained performance growth, emphasizing the need for broader and more varied question generation.
@@ 3.3 Ablation study
Figure 5: A comparison of the synthetic dataset generated by the raw instruct model, the model after QFT, the model after QPO, and the final dataset after applying reward filtering. The evaluation covers question solvability, difficulty, and instruction tuning effectiveness on Llama3-8B.
Ablation on each sub-method
@@
To validate the effectiveness of each of our sub-methods, including QFT, QPO, and reward filtering, we conducted an ablation study. We evaluated the quality of the questions generated by the models across three dimensions: solvability, difficulty, and performance in instruction tuning. To assess the model’s solvability and difficulty, we used GPT-4o-mini as the evaluation model, with the prompts provided in the Figure 11 and 12. For difficulty evaluation, we calculated the dataset’s average difficulty score based on ratings for each question: “very easy” is rated as 20 points, “easy” as 40 points, “medium” as 60 points, “hard” as 80 points, and “very hard” as 100 points.
@@
The results are shown in Figure 5. The “raw model” refers to using the instruct model to directly generate instructions and responses, as done in Xu et al. (2024). To ensure fairness, we also generated 1M question-response pairs using their method based on Qwen2-Math-7B-Instruct, which were used to train Llama3-8B. After applying QFT and QPO, the model’s performance improved across all three evaluation dimensions, demonstrating the effectiveness of our approach. Furthermore, by filtering for solvable questions and applying reward filtering to the responses, the quality of our dataset increased, resulting in significant improvements across all four evaluation benchmarks.
Question matters for data synthesis
@@
To directly compare the question quality of our constructed data with other open-source datasets, we used the same model, Qwen2-Math-7B-Instruct, to generate responses and fine-tuned DeepSeekMath-7B based on the synthetic datasets. As shown in Table 2, using the same response generation method, our model outperformed other synthetic datasets like MetaMath and OrcaMath, highlighting the high quality of our questions. NuminaMath also demonstrated competitive performance, largely due to the fact that many of its questions are drawn from real-world scenarios. This also highlights that question quality is crucial for synthetic data.
Multiple question generators enhance data diversity
We use two models as question generators: DSMath-QGen and Qwen2-Math-QGen, which are based on DeepSeekMath (Shao et al., 2024) and Qwen2-Math (Yang et al., 2024a), respectively. To explore the impact of using multiple question generators, we compared the effects of using data synthesized by a single generator versus a mix of data from both. We fixed the total dataset size at 400K and used it to fine-tune Mistral-7B. As shown in Table 3, we found that the mixed data outperformed the data generated by either single generator. A possible explanation for this improvement is the increased data diversity. In fact, we observed that DSMath-QGen tends to generate simpler, more real-world-oriented questions, while Qwen2-Math-QGen produces more challenging, theory-driven ones. From this, we recognize the potential of using multiple question generators, and we plan to incorporate more question generators as part of our future work.
Table 2: We directly compared the question quality of different open-source datasets. To ensure consistency, all responses were generated using Qwen2-Math-7B-Instruct.
Questions Source Response Synthesis Model GSM8K MATH College Math Olympiad Bench Average
MetaMath [Uncaptioned image] Qwen2-Math-7B-Instruct 84.5 53.8 40.1 22.1 50.1
OrcaMath [Uncaptioned image] Qwen2-Math-7B-Instruct 84.2 53.7 40.5 23.7 50.5
NuminaMath [Uncaptioned image] Qwen2-Math-7B-Instruct 86.0 65.9 46.1 30.2 57.1
ScaleQuest [Uncaptioned image] Qwen2-Math-7B-Instruct 89.5 66.6 47.7 29.9 58.4
@@
Table 3: The performance of Mistral-7B-v0.1 fine-tuned on ScaleQuest-DSMath, ScaleQuest-Qwen2, and a mix of both. In this setup, the instructions for ScaleQuest-DSMath and ScaleQuest-Qwen2-Math were generated by DSMath-QGen and Qwen2-Math-QGen, respectively. We fixed the training data size at 400K and found that the mixed data resulted in the greatest improvement.
Synthetic Dataset # Samples GSM8K MATH College Math Olympiad Bench Average
[Uncaptioned image] ScaleQuest-DSMath 400K 87.6 52.2 39.8 19.4 49.8
[Uncaptioned image] ScaleQuest-Qwen2-Math 400K 86.8 56.1 39.6 18.7 50.3
Mixed 400K 87.8 58.0 40.1 22.2 52.0
3.4 Cost Analysis
Table 4: Cost analysis of the entire data synthesis process. We also estimated the cost of generating the same number of tokens using proprietary models GPT-4 and GPT-4o for comparison.
@@ 3.4 Cost Analysis
The data synthesis process was conducted on a server with 8 A100-40G-PCIe GPUs. We summarize our overall costs in Table 4. Generating 1 million data samples required only 522.9 GPU hours (approximately 2.7 days on an 8-GPU server), with an estimated cost of \$680.8 for cloud server rental. [^2] https://lambdalabs.com/service/gpu-cloud This is only about 10% of the cost of generating the same data using GPT-4o. This demonstrates that our data generation method is significantly more cost-effective.
@@ 4.1 Mathematical Reasoning
4.2 Data Synthesis for Math Instruction Tuning
High-quality reasoning data, particularly well-crafted questions, is in short supply. Prior efforts have mostly started with a small set of human-annotated seed instructions and expanded them through few-shot prompting. We categorize them into two types: question-driven augmentation and knowledge-driven augmentation. Previous works focus on enhancing seed questions by introducing additional constraints or numerical changes to increase the reasoning steps required. For instance, WizardMath (Luo et al., 2023) uses a series of operations to increase the complexity of questions and answers with GPT-3.5. MetaMath (Yu et al., 2023a) enhances the questions in GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) by rewriting them in various ways, such as through semantic rephrasing, self-verification, and backward reasoning. Xwin-Math (Li et al., 2024a) and MMIQC (Liu & Yao, 2024) further explore the scalability of the synthetic data. However, these methods face a diversity challenge, as few-shot prompting often results in new instructions that are too similar to the original seed questions (Li et al., 2024b). To increase diversity, recent works have focused on knowledge-driven data synthesis, where they summarize world knowledge from the seed questions and use it to generate synthetic datasets (Didolkar et al., 2024; Shah et al., 2024). MathScale (Tang et al., 2024) extracts math concepts from seed questions and then generate math reasoning data. KPMath (Huang et al., 2024a) begins by extracting topics and key points from seed problems using a labeling model, and sample multiple topics and key points for instruction synthesis. There are other methods for enhancing dataset quality as well. DART-Math (Tong et al., 2024) focuses on enhancing the quality of responses by using rejection sampling to generate multiple correct answers for each query from GSM8K and MATH. In contrast, Numina-Math (Li et al., 2024c) improves its dataset by collecting more real-world and synthetic data, then reformatting (Fan et al., 2024) the responses using GPT-4o. This high-quality data can be integrated with our constructed dataset, resulting in an improved data mix for more effective instruction tuning.
@@ 5 Conclusion
In this work, we propose ScaleQuest, a novel data synthesis framework that unlocks the ability of open-source smaller models to independently generate large-scale, high-quality reasoning data from scratch, at a low cost. By training the problem-solving models on a small subset of questions, we effectively activate their question-generation capabilities. We also introduce a response enhancement method. With these techniques, we successfully developed a fully synthetic math reasoning dataset consisting of 1 million question-answer pairs. Using this dataset, we fine-tuned the model and achieved remarkable improvements, with gains ranging from 29.2% to 46.4% compared to the base model. The fine-tuned 7B model, Qwen2-Math-7B-ScaleQuest, outperforms all competitors in the 7B-70B range and even surpasses proprietary models like GPT-4-Turbo and Claude-3.5-Sonnet.
Due to time and cost constraints, there are several areas where our approach can be further optimized. For instance, leveraging more powerful, larger problem-solving models like Qwen2.5-Math-72B-Instruct (Yang et al., 2024b) for question and response generation, using advanced models such as GPT-4o for constructing preference data for Question Preference Optimization, and further scaling up the generation of synthetic data. Each stage of our process has significant room for improvement. In this paper, we have demonstrated the potential of this framework, laying the groundwork for future enhancements.
Furthermore, despite the progress made in this work, there are still several limitations that need to be addressed. In our future research, we will concentrate on the following areas:
  • Large-scale and diverse high-quality data: This work chooses mathematical reasoning as a case study to demonstrate the effectiveness of our method. In the future, we will focus on broader and more complex tasks such as science and competitive programming. Additionally, future research will aim to continuously scale data synthesis to explore the scaling laws for synthetic data and seek a more efficient approach to scaling data generation.
  • Self-improvement capability: Our experiments demonstrate the model’s self-improvement capability, meaning that it can generate data of higher quality than its original training set. This is evident as Qwen2-Math-7B-ScaleQuest slightly outperforms Qwen2-Math-7B-Instruct. To further explore the upper bounds of self-improvement, our future research will focus on synthesizing preference-tuning data to better align the LLMs.
Appendix A Additional Data Statistics
@@
We analyze the dataset coverage through two aspects: (1) Problem Topic Coverage, such as algebra and geometry. Following Huang et al. (2024a), we use GPT-4o to categorize the topics of the given questions, with prompt illustrated in Figure 13. Figure 8 presents the results. We found that the topics covered the major areas of mathematics, such as arithmetic, algebra, geometry, and others. (2) Embedding space analysis. Following Zhao et al. (2024) and Xu et al. (2024), we first compute the input embeddings of the questions and then project them into a two-dimensional space using t-SNE (Van der Maaten & Hinton, 2008). We included only real-world datasets, such as GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and NuminaMath (Li et al., 2024c) (which contains a small portion of synthetic questions). As shown in Figure 8, our synthetic data closely resembles the real-world questions.
@@
We used Llama3-8B-Guard (Inan et al., 2023) as a discriminator model to detect any unsafe elements in the data. After sampling 10K instances from the 1 million samples, we found that only 0.1% were flagged as unsafe.
@@
We sampled several generated examples from our datasets, as shown in Figure 16, 17 and 18. The generated math problems are of high quality, driving effective learning.
@@
MMIQC (Liu & Yao, 2024) 2294K GPT-4 & GPT-3.5-Turbo & Human

### v2

@@ 3.2 Main Results
ScaleQuest significantly outperforms other baselines
Table 1 presents the results. ScaleQuest significantly outperforms previous synthetic methods, with average performance improvements ranging from 5.6% to 11.5% over the prior state-of-the-art (SoTA) on both general base models and math-specialized foundation models. Qwen2-Math-7B-ScaleQuest achieve a zero-shot pass@1 accuracy of 73.4 on the MATH benchmark, matching the performance of GPT-4-Turbo. For out-of-domain tasks, Qwen2-Math-7B-ScaleQuest outperform its teacher model, Qwen2-Math-7B-Instruct, with scores of 89.7 on the GSM8K benchmark, 73.4 on the MATH benchmark, and 38.5 on the Olympiad benchmark. In this experiment, we do not strictly control for identical training data volumes due to practical constraints (i.e., some datasets are not publicly available). To address this, we compare the results with publicly available datasets under equal data volumes in Appendix E.2. It’s important to highlight that Qwen2-Math-7B-Instruct has undergone preference alignment, utilizing the powerful reward model Qwen2-Math-RM-72B (Yang et al., 2024a), while our model is only an instruction tuning version.
@@
We also compare the scalability of our dataset with other publicly available datasets, including MetaMath (Yu et al., 2023a), DART-Math (Tong et al., 2024), and Numina-Math (Li et al., 2024c). We train Llama3-8B on these datasets and compare their performance changes with increased data size. The results are presented in Figure 1. For the in-domain evaluation (MATH), our method demonstrates high data efficiency, achieving superior results with the same amount of data. In out-of-domain evaluations (Olympiad Bench), it also shows strong scalability, continuing to improve even as other datasets reach their limits. A limited question set leads to constrained improvements in model performance, as demonstrated by the results of DART-Math, which relies on a small number of questions and generates numerous correct answers through rejection sampling. Our results further demonstrate that diverse questions support sustained performance growth, emphasizing the need for broader and more varied question generation.
@@ 3.3 Ablation study
Figure 5: A comparison of the synthetic dataset generated by the raw instruct model, the model after QFT, the model after QPO, and the final dataset after applying reward filtering. Left: The solvable ratio and difficulty score of the generated questions. The solvable ratio refers to the proportion of generated questions that are judged as “solvable”, while the difficulty score represents the average difficulty rating assigned to each generated question. For difficulty evaluation, we calculate the dataset’s average difficulty score based on ratings for each question: “very easy” is rated as 20 points, “easy” as 40 points, “medium” as 60 points, “hard” as 80 points, and “very hard” as 100 points. Right: The instruction tuning effectiveness on Llama3-8B-Base.
Each sub-method contributes to the final performance
@@
To validate the effectiveness of each of our sub-methods, including QFT, QPO, and reward filtering, we conduct an ablation study. We evaluate the quality of the questions generated by the models across three dimensions: solvability, difficulty, and performance in instruction tuning. For solvability and difficulty, we use the same evaluation process as stated in section 2.3.
@@
The results are shown in Figure 5. The “raw model” refers to using the instruct model to directly generate instructions and responses, as done in Xu et al. (2024). To ensure fairness, we also generate 1M question-response pairs using their method based on Qwen2-Math-7B-Instruct, which are used to train Llama3-8B-Base. After applying QFT and QPO, the model’s performance improves across all three evaluation dimensions, demonstrating the effectiveness of each sub-method.
Question Source GSM8K MATH CM OB Avg
MetaMath 84.5 53.8 40.1 22.1 50.1
OrcaMath 84.2 53.7 40.5 23.7 50.5
NuminaMath 86.0 65.9 46.1 30.2 57.1
ScaleQuest 89.5 66.6 47.7 29.9 58.4
Table 2: Comparison of question quality across different datasets. To ensure consistency, all responses are generated using Qwen2-Math-7B-Instruct with the same reward filtering process. CM and OB refer to College Math and Olympiad Bench, respectively. Additional results based on the other three backbone models are shown in Appendix E.4.
Comparison of question quality
@@
To directly compare the question quality of our constructed data with other open-source datasets, we use the same model, Qwen2-Math-7B-Instruct, to generate responses. We then fine-tune DeepSeekMath-7B-Base using these synthetic datasets. As shown in Table 2, our model outperforms other synthetic datasets like MetaMath and OrcaMath, highlighting the high quality of our questions. NuminaMath also demonstrates competitive performance, largely due to the fact that many of its questions are drawn from real-world scenarios. This highlights the importance of question quality for synthetic data.
Dataset GSM8K MATH CM OB Avg
[Uncaptioned image] SQ-DSMath 87.6 52.2 39.8 19.4 49.8
[Uncaptioned image] SQ-Qwen2 86.8 56.1 39.6 18.7 50.3
Mixed 87.8 58.0 40.1 22.2 52.0
@@
Table 3: The performance of Mistral-7B fine-tuned on ScaleQuest-DSMath, ScaleQuest-Qwen2, and a mix of both. We fix the training data size at 400K and find that the mixed data results in the largest improvement.
@@
Table 4: Cost analysis of the entire data synthesis process. We also estimate the cost of generating the same number of tokens using proprietary models GPT-4 and GPT-4o for comparison.
Multiple question generators enhance data diversity
We use two models as question generators: DSMath-QGen and Qwen2-Math-QGen, which are based on DeepSeekMath (Shao et al., 2024) and Qwen2-Math (Yang et al., 2024a), respectively. To explore the impact of using multiple question generators, we compare the effects of using data synthesized by a single generator versus a mix of data from both. As shown in Table 3, we find that the mixed data outperforms the data generated by either single generator. A possible explanation for this improvement is the increased data diversity. In fact, we observe that DSMath-QGen tends to generate simpler, more real-world-oriented questions, while Qwen2-Math-QGen produces more challenging, theory-driven ones.
3.4 Human Evaluation Results
We conduct a human evaluation of the generated data, focusing on three aspects: clarity, reasonableness, and real-world relevance. For reference, we also include two high-quality, human-curated datasets, GSM8K and MATH. A total of 40 examples are sampled from each dataset and evaluated based on clarity, coherence, and real-world relevance, with scores ranging from 1 to 5. The results are presented in Table 5. In terms of clarity and reasonableness, our synthetic data surpasses NuminaMath but still falls short of the high-quality, real-world datasets like the training sets of GSM8K and MATH. Regarding real-world relevance, GSM8K leans toward practical, real-life scenarios, while MATH focuses more on theoretical mathematical derivations. Our generated data can be seen as a balance between the two.
Dataset Clarity Reasonableness Real-world relevance
GSM8K 4.4 4.5 3.9
MATH 4.1 4.3 2.4
NuminaMath 3.8 4.0 2.4
ScaleQuest 3.9 4.0 2.8
Table 5: Human Evaluation Results.
3.5 Cost Analysis
The data synthesis process is conducted on a server with 8 A100-40G-PCIe GPUs. We summarize our overall costs in Table 4. Generating 1 million data samples requires only 522.9 GPU hours (approximately 2.7 days on an 8-GPU server), with an estimated cost of \$680.8 for cloud server rental. [^2] https://lambdalabs.com/service/gpu-cloud This is only about 10% of the cost of generating the same data using GPT-4o, demonstrating the cost-effectiveness of our method.
@@ 4.1 Mathematical Reasoning
4.2 Data Synthesis for Instruction Tuning
High-quality reasoning data, particularly well-crafted questions, is in short supply. Prior efforts have mostly started with a small set of human-annotated seed instructions and expanded them through few-shot prompting. We categorize them into two types: question-driven augmentation (Luo et al., 2023; Yu et al., 2023a; Li et al., 2024a; Liu and Yao, 2024; Li et al., 2024b) and knowledge-driven augmentation (Didolkar et al., 2024; Shah et al., 2024; Tang et al., 2024; Huang et al., 2024a). There are other methods for enhancing dataset quality as well. DART-Math (Tong et al., 2024) focuses on enhancing the quality of responses by using difficulty-aware rejection sampling. In contrast, Numina-Math (Li et al., 2024c) improves its dataset by collecting more real-world and synthetic data. These high-quality datasets can be integrated with our constructed dataset, resulting in an improved data mix for more effective instruction tuning.
@@ 5 Conclusion
In this work, we propose ScaleQuest, a novel data synthesis framework that unlocks the ability of lightweight models to independently generate large-scale, high-quality reasoning data from scratch, at a low cost. Using this dataset, we fine-tune the model and achieve remarkable improvements, with gains ranging from 29.2% to 46.4% compared to the base model, outperforming the strong baselines.
Limitation
We explore the potential of lightweight open-sourced models to generate high-quality instruction tuning data. However, the effectiveness on larger and more powerful models, such as Qwen2.5-Math-72B-Instruct and Llama3.3-70B-Instruct, remains uncertain. Our future research will focus on experimenting with these larger models. Additionally, although the optimizations for solvability and difficulty have shown positive effects, the model-generated responses still fall short of being fully satisfactory. Therefore, further improvements in question preference alignment are crucial and will be an important direction for our future work.
Acknowledgements
We want to thank all the anonymous reviewers for their valuable comments. This work was supported by the National Science Foundation of China (NSFC No. 62206194), the Natural Science Foundation of Jiangsu Province, China (Grant No. BK20220488), the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Priority Academic Program Development of Jiangsu Higher Education Institutions, and Sponsored by CCF-Tencent Rhino-Bird Open Research Fund. We also acknowledge MetaStone Tech. Co. for providing us with the software, optimisation on high performance computing and computational resources required by this work.
Appendix A Hyper-parameters settings
Training Problem Designers
During the QFT stage, both models are trained on a mixed training subset of GSM8K and MATH problems, containing a total of 15K problems. We train for only 1 epoch, considering that training for more epochs might cause the models to overfit the training problems and negatively impact the diversity of generated questions. We also use sequence packing (Krell et al., 2021) to accelerate training. In the QPO stage, we use 10K preference data for training, with a learning rate of 5e-7 and a batch size of 128.
Question Generation
During this process, we set the maximum generation length to 512, a temperature of 1.0, and a top-p value of 0.99. To ensure quality, we apply a question filtering pipeline (section 2.4) that involves language filtering, solvability filtering, and difficulty sampling. This process refines the dataset, leaving approximately 1M questions, 400K from DeepSeek-QGen and 600K from Qwen2-Math-QGen.
Response Generation
In the process, we set the maximum generation length to 2048, with a temperature of 0.7 and top-p of 0.95. We use chain-of-thought prompt (Wei et al., 2022) to synthesize solutions. We use vLLM (Kwon et al., 2023) to accelerate the generation and Ray (Moritz et al., 2018) to deploy distributed inference.
Instruction Tuning
All models are fine-tuned for 3 epochs in our experiments unless specified otherwise. We use a linear learning rate schedule with a 3% warm-up ratio, reaching a peak of 5e-5 for Llama3 and DeepSeekMath and 1e-5 for the other models, followed by cosine decay to zero.
Appendix B Additional Data Statistics
@@
We analyze the dataset coverage through two aspects: (1) Problem Topic Coverage, such as algebra and geometry. Following Huang et al. (2024a), we use GPT-4o to categorize the topics of the given questions, with prompt illustrated in Figure 14. Figure 8 presents the results. We find that the topics cover the major areas of mathematics, such as arithmetic, algebra, geometry, and others. (2) Embedding space analysis. Following Zhao et al. (2024) and Xu et al. (2024), we first compute the input embeddings of the questions and then project them into a two-dimensional space using t-SNE (Van der Maaten and Hinton, 2008). We include only real-world datasets, such as GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and NuminaMath (Li et al., 2024c) (which contains a small portion of synthetic questions). As shown in Figure 8, our synthetic data closely resembles the real-world questions.
@@
Data Leakage Analysis
We conduct an n-gram similarity analysis between the generated questions and all test sets from both our dataset and other baseline datasets. Based on prior empirical analysis (Brown, 2020; Wei et al., 2021), we set n=13 to prevent spurious collisions and calculate how much the test sets overlap with training data to assess data contamination. Table 6 shows the clean ratio of our dataset and other baseline datasets. The results demonstrate that our dataset achieves a relatively high level of data cleanliness compared to other datasets, suggesting that our method generates novel questions instead of memorizing existing ones.
Dataset GSM8K MATH CM OB Avg
MetaMath 99.8 92.2 100 99.7 97.9
NuminaMath 99.8 89.8 99.9 86.8 94.1
DART-Math 99.8 91.5 100.0 99.6 97.7
MMIQC 99.8 88.0 98.9 97.9 96.2
SQ (Ours) 99.9 92.8 99.8 97.2 97.4
Table 6: Overlap statistics for the datasets used. We report the clean ratio of the test set, representing the percentage of test samples that have no matching n-grams with samples in the training set.
@@
We use Llama3-8B-Guard (Inan et al., 2023) as a discriminator model to detect any unsafe elements in the data. After sampling 10K instances from the 1 million samples, we find that only 0.1% are flagged as unsafe.
@@
We sample several generated examples from our datasets, as shown in Figure 17, 18 and 19. The generated math problems are of high quality, driving effective learning.
@@
MMIQC (Liu and Yao, 2024) 2294K GPT-4 & GPT-3.5-Turbo & Human